OpenStudy (anonymous):

Simplify 9y^7/15x^5•5x^2/45y^3

OpenStudy (shamim):

tell me y^3/y=?

OpenStudy (anonymous):

We know: \(\huge \color{red}{\frac{a^x}{a^y}=a^{x-y}}\) \(\Large \frac{9y^7}{15x^5}\times \frac{5x^2}{45y^3}=\frac{9y^7}{45y^3}\times \frac{5x^2}{15x^5}\) Now, tell me \(\Large \frac{9y^7}{45y^3}=?\) and \(\Large \frac{5x^2}{15x^5}=?\)

OpenStudy (anonymous):

\(\Large \frac{9y^7}{45y^3}=\frac{9}{45}\times \frac{y^7}{y^3}=\frac{1}{5}\times y^{7-3}=\frac{y^4}{5}\) \(\Large \frac{5x^2}{15x^5}=\frac{5}{15}\times \frac{x^2}{x^5}=\frac{1}{3}\times x^{2-5}=\frac{x^{-3}}{3}=\frac{1}{3x^3}\) \(\Large \color{red}{(a^{-m}=\frac{1}{a^m} \Rightarrow x^{-3}=\frac{1}{x^3})}\)